in an exeriment, 4 gram of M2Ox oxide was reduced to 2.8 gram of the metal.if the atomic mass of the metal is 56 gram per mole,the number of o-atoms in the oxide is

Grade:12th pass

1 Answers

Vikas TU
7 years ago
moles of M2Ox => 4/(2M + 16x)
and moles of M => 8/(2M + 16x)
From 2.8 gm,
moles of M => 2.8/M
Now both should be equal hence,
 2.8/M = 4/(M + 8x)
0.7M = 1/(M + 8x)............(1)
Now given M = 56 put the value in eqn. (1) we get,
39.2 = 1/(8x + 56)
(x + 7) = 1/313.6
 x = 7 (approx) cant be negative.